AshCar82:

Compute the amount of interest earned in the following simple interest problem. A deposit of $5,000 at 8.5% for 120 days = _____. $1,402.50 $141.67 $14,025 $51,000

mikewwe13:

To compute the amount of interest earned in this simple interest problem, we can use the formula: I = P * r * t Where: I = interest earned P = principal amount (deposit) r = interest rate (as a decimal) t = time (in years) First, we need to convert the time period of 120 days to years by dividing by the number of days in a year (365): t = 120/365 t = 0.3288 years Next, we can plug in the values we have into the formula and solve for I: I = 5,000 * 0.085 * 0.3288 I = $141.67 Therefore, the amount of interest earned in this simple interest problem is $141.67.